# VeOmni: Scaling Any Modality Model Training with Model-Centric Distributed Recipe Zoo

### v2

@@ 4.1 Experimental Setup
We evaluate the training performance and scalability of VeOmni on large-scale productive GPU clusters across configurations ranging from 8 to 128 GPUs, enabling a comprehensive study of VeOmni’s behavior under both moderate and large-scale distributed settings.
Models and Datasets.
@@
We evaluate a diverse set of model architectures, including dense models such as Qwen2-VL 7B and 72B [65], and a mixture-of-experts (MoE) omni-modal LLM based on Qwen3-MoE [57]. To assess the multimodal capabilities of VeOmni, we use the following domain-specific datasets: FineWeb-100T [43] for text understanding, ShareGPT4V [5] for image understanding, LLaVA-Video [76] for video understanding, Voice Assistant [69] for audio understanding, and ImageNet [13] for image generation tasks. Each model is adapted using its native instruction template and augmented with special tokens to delineate modality boundaries (e.g. <image_start> and <image_end>).
Workloads and Metrics.
@@
To assess scalability across different model sizes and GPU counts, we progressively scale the input context length from 8K to 256K tokens. We evaluate the performance and scalability of VeOmni by measuring training throughput (tokens per second) and Memory Footprint Utilization (MFU) [7], which together provide a comprehensive view of system efficiency under varying workloads and parallel strategies. Additionally, we analyze loss convergence behavior across three structurally distinct omni-modal LLMs to validate training stability and overall effectiveness. For all experiments, we freeze the modality-specific encoders and decoders, while fully fine-tuning the remaining components, including the foundation model and multimodal projectors.
4.2 Training Recipes under Different Scenarios
@@
Figure 4: 2D Parallelism (FSDP+SP) on Qwen2-VL 7B with 8 GPUs. The maximum sequence length varies from 8K to 256K tokens, with supported sequence parallel sizes ranging from 1 to 4.
@@
Figure 5: 2D Parallelism (FSDP+SP) on Qwen2-VL 72B with 128 GPUs. The maximum sequence length varies from 8K to 96K tokens, with supported sequence parallel sizes ranging from 1 to 8.
@@
Figures 46 (a) present a comprehensive comparison of the efficiency of various parallelism strategies—fully sharded data parallel (FSDP), sequence parallelism (SP), and expert parallelism (EP)—across diverse model scales and hardware configurations. Figures 4-5 show results of training Qwen2-VL on 8 and 128 GPUs, respectively. The findings reveal that increasing the degree of sequence parallelism enables the models to handle significantly longer context lengths. In particular, VeOmni can support up to 192K tokens for training the 7B model with an MFU of 61.5% and 96K tokens for training the 72B model with an MFU of 54.82%. Figure 6 further extends this analysis to a 3D parallelism configuration (FSDP + SP + EP) using a 30B parameter a mixture-of-experts (MoE) omni-modal LLM based on Qwen3-MoE, on 128 GPUs. In this setting, combinations involving moderate levels of SP and EP achieve a balanced trade-off, supporting extended sequence lengths of up to 160K tokens, while maintaining competitive throughput.
@@
The experimental results underscore the efficiency of VeOmni in handling long-sequence training and scaling mixture-of-experts (MoE) models. Specifically, VeOmni demonstrates strong performance by minimizing overhead in short-context scenarios, while effectively leveraging sequence and expert parallelism to maintain scalability and feasibility in extended-context and MoE settings.
@@ 4.3 Convergence Study on Omni-Modal LLMs
Figure 6: (a) 3D Parallelism (FSDP+SP+EP) on a 30B Qwen3MoE-based omni-modal LLM with 128 GPUs. The maximum sequence length varies from 16K to 192K tokens, with supported sequence parallel sizes ranging from 1 to 4 and expert parallel sizes ranging from 1 to 4. (b) Convergence Study on three Distinct omni-modal LLMs.
Figure 6 (b) presents the convergence behavior of three omni-modal LLMs on multimodal understanding and generation tasks. The understanding tasks span four modalities, i.e., text, image, video, and audio, while the generation tasks involve text and image synthesis. Janus [67] is trained solely on image understanding and image generation. LLaMA#Omni and Qwen3-Moe#Omni are two custom models that share similar architectures: Qwen2.5-Omni [70] ViT serves as the image and video encoder, Qwen2.5-Omni Whisper as the audio encoder, and MoVQGAN [79] as the image decoder. LLaMA#Omni uses LLaMA [60] as the foundation, while Qwen3-Moe#Omni adopts Qwen3-MoE [57]. Janus adopts SigLip [73] as the image encoder, LLaMA as the foundation model, and LlamaGen [53] as the image decoder. “LM loss” refers to the cross-entropy loss over text tokens, and “decoder loss” refers to the cross-entropy loss over image tokens. As shown in Figure 6 (b), all models exhibit stable convergence across both understanding and generation tasks, demonstrating that VeOmni enables efficient and robust training for large omni-modal LLMs.
@@ Acknowledgments
Appendix A Technical Appendices and Supplementary Material
A.1 Controlled Benchmarking of VeOmni vs. TorchTitan on Large Language Models
In the main paper, we demonstrate that VeOmni supports efficient large-scale training for multimodal models, a capability not yet supported by other existing frameworks. As such, a direct comparison in multimodal training scenarios is infeasible. To provide a fair benchmark, we conduct controlled experiments on large-scale text-only models to compare VeOmni with TorchTitan [36], a state-of-the-art distributed training framework.
Table 1: Performance Comparison between VeOmni and TorchTitan on Qwen2-7B, 128GPUs. Both use mixed precision training and full activation checkpointing. micro batch size of 1, global batch size of 128.
Techniques Sequence Length Memory Throughput MFU
@@ Appendix A Technical Appendices and Supplementary Material
FSDP 8k 21.91GB 23.26GB 6940 6639 38.76 34.61
FSDP+SP4 8k 23.23GB 22.55GB 3830 3590 20.66 18.72
FSDP 16k 24.45GB 39.58GB 6583 6082 43.22 37.62
FSDP+SP4 16k 23.92GB 25.08GB 4889 4056 31.92 25.24
FSDP 32k 31.41GB 71.81GB 5315 4878 44.42 39.68
FSDP+SP4 32k 25.27GB 30.13GB 4629 3727 38.69 30.32
FSDP 64k 43.00GB OOM 3725 OOM 45.92 OOM
FSDP+SP4 64k 28.70GB 44.82GB 3452 2853 42.35 34.32
FSDP 128k 70.26GB OOM 2258 OOM 44.95 OOM
FSDP+SP4 128k 35.49GB OOM 2187 OOM 43.49 OOM
Table 2: Performance Comparison between VeOmni and TorchTitan on Qwen2.5-32B, 128GPUs. Both use mixed precision training and full activation checkpointing. Micro batch size of 1, global batch size of 128.
Techniques Sequence Length Memory Throughput MFU
@@
FSDP+SP4 8k 35.98GB 34.37GB 1140 1017 26.05 22.90
FSDP+SP8 8k 36.47GB 33.42GB 668 596 15.14 13.43
FSDP+SP4 16k 37.43GB 39.48GB 1316 1131 34.52 29.05
FSDP+SP8 16k 37.31GB 37.14GB 1074 789 27.87 20.28
FSDP+SP4 32k 42.42GB 46.78GB 1240 1022 36.11 32.75
FSDP+SP8 32k 38.76GB 44.04GB 1115 821 40.36 26.31
FSDP+SP4 64k 49.06GB 68.16GB 957 817 43.02 36.58
FSDP+SP8 64k 43.75GB 58.58GB 903 689 40.74 30.84
FSDP+SP4 128k 64.85GB OOM 632 OOM 44.20 OOM
FSDP+SP8 128k 50.41GB 79.64GB 613 507 42.92 35.57
Table 3: Performance Comparison between VeOmni and TorchTitan on Qwen2-72B, 128GPUs. Both use mixed precision training and full activation checkpointing. Micro batch size of 1, global batch size of 128.
Techniques Sequence Length Memory Throughput MFU
@@
FSDP+SP4 8k 58.62GB 54.08GB 565 499 28.04 24.74
FSDP+SP8 8k 58.35GB 51.73GB 330 297 16.11 14.75
FSDP+SP4 16k 62.15GB 64.35GB 624 544 35.34 30.44
FSDP+SP8 16k 59.98GB 58.02GB 532 394 29.82 22.03
FSDP+SP4 32k 68.35GB 79.36GB 592 497 41.05 34.16
FSDP+SP8 32k 63.49GB 70.39GB 533 398 36.72 27.31
FSDP+SP4 64k 79.10GB OOM 468 OOM 43.98 OOM
FSDP+SP8 64k 70.09GB 79.04GB 441 335 41.63 31.56
FSDP+SP4 128k OOM OOM OOM OOM OOM OOM
FSDP+SP8 128k 76.67GB OOM 303 OOM 43.68 OOM
Table 4: Performance of VeOmni on Qwen3Moe-30B, 128GPUs. Use mixed precision training and full activation checkpointing. Micro batch size of 1, global batch size of 128. (TorchTitan not supported)
Techniques Sequence Length Memory Throughput MFU
FSDP+SP1+EP8 8k 45.58GB 2776 7.45
FSDP+SP4+EP8 8k 22.56GB 1216 3.02
FSDP+SP1+EP8 16k 69.99GB 2853 10.74
FSDP+SP4+EP8 16k 34.32GB 1897 6.55
FSDP+SP1+EP8 32k 77.25GB 2417 13.40
FSDP+SP4+EP8 32k 63.08GB 2064 11.35
FSDP+SP1+EP8 64k OOM OOM OOM
FSDP+SP4+EP8 64k 77.92GB 1716 15.65
FSDP+SP1+EP8 128k OOM OOM OOM
FSDP+SP4+EP8 128k 78.01GB 1075 17.92
These benchmark results demonstrate that VeOmni consistently achieves higher throughput and memory efficiency across a range of model sizes and sequence lengths. Notably, VeOmni supports large-sequence and MoE models that exceed the memory or capability limits of TorchTitan.
Appendix B API Design
B.1 OmniData Process and Collator API
Listing LABEL:lst:veomni-data-collator demonstrates the data collation pipeline in VeOmni, which provides a unified interface for processing and batching multimodal inputs, including text, image, video, and audio. The OmniDataCollatorWithPacking class enables fine-grained control over feature packing and concatenation, making it straightforward to prepare batched training data for diverse modalities. This modular design simplifies the integration of new modalities and supports flexible token-level and feature-level alignment strategies. The resulting train_dataloader is constructed using build_dataloader(), which automatically aligns with the parallel training configuration.
Listing 1: VeOmni OmniDataCollator API
@@ B.1 OmniData Process and Collator API
OmniDataCollatorWithPacking,
build_dataloader,
@@
B.2 Building Omni-Modal LLMs
Listing LABEL:lst:veomni-build-omni-model shows the usage of VeOmni’s build_omni_model API, which serves as the unified entry point for constructing multimodal OmniModels. Given the configuration file and model component paths (including modality-specific encoders, decoders, and the foundation model), this function automatically assembles and initializes the model on a designated device. By abstracting away the underlying initialization logic, VeOmni ensures that users can easily instantiate complex multimodal models in a modular and scalable fashion, while maintaining compatibility with distributed parallelization workflows.
Listing 2: build_omni_model API usage
@@ B.2 Building Omni-Modal LLMs
model: SeedOmniModel = build_omni_model(
@@
B.3 Forward Process of OmniModel in VeOmni
Listing LABEL:lst:veomni-train-omni-model-LABEL:lst:veomni-infer-omni-model illustrates the partial forward process implementation of the plug-and-play architecture in VeOmni, named OmniModel, which adopts an encoder–foundation–decoder architecture. The snippet demonstrates how the forward logic across training and inference modes integrates each module—encoder, foundation model, and decoder—in a seamless and modular pipeline. During inference, the control flow dynamically switches to modality-specific generation branches based on the generated text special tokens, enabling flexible multimodal outputs. During inference, OmniModel supports parallelized generation and classifier-free guidance [67]. This architecture highlights the clear decoupling between encoder, foundation, and decoder modules, allowing each component to define and execute its own forward logic independently.
Listing 3: Training Process
@@ B.3 Forward Process of OmniModel in VeOmni
Listing 4: Inference Process
class OmniModel():
@@
B.4 Parallel State of VeOmni
Listing LABEL:lst:veomni-parallel-state demonstrates the core interface of VeOmni’s parallel_state API, which provides a unified and declarative initialization of all N-dimensional parallel configurations. Unlike traditional process group–based parallel management, VeOmni leverages a DeviceMesh abstraction to represent and organize various parallel topologies, including data parallel (DP), tensor parallel (TP), expert parallel (EP), pipeline parallel ($P$), and Ulysses-based sequence parallel (SP). Once initialized via init_parallel_state, users can conveniently access mesh and group objects, ranks, and communication topologies through the global parallel_state object. This design simplifies the orchestration of hybrid-parallel strategies and decouples low-level distributed logic from model implementation.
Listing 5: parallel_state API
from veomni.distributed.parallel_state import get_parallel_state, init_parallel_state

### v3

@@ 4.1 Experimental Setup
Environments. We evaluate the training performance and scalability of VeOmni on large-scale productive GPU clusters across configurations ranging from 8 to 128 GPUs, enabling a comprehensive study of VeOmni’s behavior under both moderate and large-scale distributed settings.
@@
Models and Datasets. We evaluate a diverse set of model architectures, including dense models such as Qwen2-VL 7B and 72B [66], and a mixture-of-experts (MoE) omni-modal LLM based on Qwen3-30B-A3B [58]. To assess the multimodal capabilities of VeOmni, we use the following domain specific datasets: FineWeb-100T [43] for text understanding, ShareGPT4V [5] for image understanding, LLaVA-Video [78] for video understanding, Voice Assistant [70] for audio understanding, and ImageNet [13] for image generation tasks. Each model is adapted using its native instruction template and augmented with special tokens to delineate modality boundaries (e.g. <image_start> and <image_end>).
@@
Workloads and Metrics. To assess scalability across different model sizes and GPU counts, we progressively scale the input context length from 8K to 256K tokens. We evaluate the performance and scalability of VeOmni by measuring training throughput (tokens per second per GPU) and model FLOPs utilization (MFU) [7], which together provide a comprehensive view of system efficiency under varying workloads and parallel strategies. Additionally, we analyze loss convergence behavior across three structurally distinct omni-modal LLMs to validate training stability and overall effectiveness. For all experiments, we freeze the modality-specific encoders and decoders, while fully fine-tuning the remaining components, including the foundation model and multimodal projectors.
4.2 Comparison of Training Recipes under Different Scenarios
@@
Figure 4: 2D Parallelism (FSDP+SP) with Qwen2-VL [66] 7B on 8 GPUs. The maximum context length varies from 8K to 256K, with allowed sequence parallel sizes ranging from 1 to 4.
@@
Figure 5: 2D Parallelism (FSDP+SP) with Qwen2-VL [66] 72B on 128 GPUs. The maximum context length varies from 8K to 96K, with allowed sequence parallel sizes ranging from 1 to 8.
@@
Figures 46 (a) present a comprehensive comparison of the efficiency of various training recipes across different model sizes and GPU numbers. Figures 4-5 show results of training Qwen2-VL [66] on 8 and 128 GPUs, respectively. The findings reveal that increasing the degree of sequence parallelism enables the models to handle significantly longer context lengths. In particular, VeOmni can support up to 192K context length for training the 7B model with an MFU of 61.5% and 96K context length for training the 72B model with an MFU of 54.82%. Figure 6 further extends this analysis to a 3D parallelism configuration (FSDP+SP+EP) using a 30B parameter a mixture-of-experts (MoE) omni-modal LLM based on Qwen3-30B-A3B [58], on 128 GPUs. In this setting, combinations involving moderate levels of SP and EP achieve a balanced trade-off, supporting extended context lengths of up to 160K, while maintaining competitive throughput.
@@
The experimental results underscore the efficiency of VeOmni in handling long-sequence training and scaling mixture-of-experts (MoE) models. Specifically, VeOmni demonstrates strong performance by minimizing overhead in short sequence scenarios, while effectively leveraging sequence and expert parallelism to maintain scalability and feasibility in long sequence and MoE settings.
@@ 4.3 Convergence Study on Omni-Modal LLMs
Figure 6: (a) 3D parallelism (FSDP+SP+EP) on a 30B MoE-based omni-modal LLM with 128 GPUs. The maximum sequence length varies from 16K to 192K, with allowed sequence parallel sizes ranging from 1 to 4 and expert parallel sizes ranging from 1 to 4. (b) Convergence study on three distinct omni-modal LLMs.
Figure 6 (b) presents the convergence behavior of three omni-modal LLMs on multimodal understanding and generation tasks. The understanding tasks span four modalities, i.e., text, image, video, and audio, while the generation tasks involve text and image synthesis. Janus [68] is trained solely on image understanding and image generation. LLaMA#Omni and Qwen3-MoE#Omni are two custom models that share similar architectures: Qwen2.5-Omni [71] ViT serves as the image and video encoder, Qwen2.5-Omni Whisper as the audio encoder, and MoVQGAN [81] as the image decoder. LLaMA#Omni uses LLaMA [61] as the foundation model, while Qwen3-MoE#Omni adopts Qwen3-30B-A3B [58]. Janus adopts SigLip [75] as the image encoder, LLaMA as the foundation model, and LlamaGen [54] as the image decoder. “LM loss” refers to the cross-entropy loss over text tokens, and “decoder loss” refers to the cross-entropy loss over image tokens. As shown in Figure 6 (b), all models exhibit stable convergence across both understanding and generation tasks, demonstrating that VeOmni enables efficient and robust training for large omni-modal LLMs.
@@ Acknowledgments
Appendix A Comparison of VeOmni and TorchTitan on Training Large Language Models
In the main paper, we demonstrate that VeOmni supports efficient large-scale training for omni-modal models, a capability not yet supported by other existing frameworks. As such, a direct comparison in omni-modal training scenarios is infeasible. To provide a fair benchmark, we conduct controlled experiments on large-scale text-only models to compare VeOmni with TorchTitan [36], a state-of-the-art distributed training framework.
Table 1: Performance comparison between VeOmni and TorchTitan with Qwen2-7B [66] on 128GPUs. Both configurations adopt mixed-precision training and full activation checkpointing, with a micro-batch size of 1 and a global batch size of 128.
Method Seqlen Memory (GB) Throughput MFU (%)
@@ Appendix A Comparison of VeOmni and TorchTitan on Training Large Language Models
FSDP 8k 21.91 23.26 6,940 6,639 38.76 34.61
FSDP+SP4 8k 23.23 22.55 3,830 3,590 20.66 18.72
FSDP 16k 24.45 39.58 6,583 6,082 43.22 37.62
FSDP+SP4 16k 23.92 25.08 4,889 4,056 31.92 25.24
FSDP 32k 31.41 71.81 5,315 4,878 44.42 39.68
FSDP+SP4 32k 25.27 30.13 4,629 3,727 38.69 30.32
FSDP 64k 43.00 OOM 3,725 - 45.92 -
FSDP+SP4 64k 28.70 44.82 3,452 2,853 42.35 34.32
FSDP 128k 70.26 OOM 2,258 - 44.95 -
FSDP+SP4 128k 35.49 OOM 2,187 - 43.49 -
Table 2: Performance comparison between VeOmni and TorchTitan with Qwen2.5-32B [45] on 128GPUs. Both configurations adopt mixed-precision training and full activation checkpointing, with a micro-batch size of 1 and a global batch size of 128.
Method Seqlen Memory (GB) Throughput MFU (%)
@@
FSDP+SP4 8k 35.98 34.37 1,140 1,017 26.05 22.90
FSDP+SP8 8k 36.47 33.42 668 596 15.14 13.43
FSDP+SP4 16k 37.43 39.48 1,316 1,131 34.52 29.05
FSDP+SP8 16k 37.31 37.14 1,074 789 27.87 20.28
FSDP+SP4 32k 42.42 46.78 1,240 1,022 36.11 32.75
FSDP+SP8 32k 38.76 44.04 1,115 821 40.36 26.31
FSDP+SP4 64k 49.06 68.16 957 817 43.02 36.58
FSDP+SP8 64k 43.75 58.58 903 689 40.74 30.84
FSDP+SP4 128k 64.85 OOM 632 - 44.20 -
FSDP+SP8 128k 50.41 79.64 613 507 42.92 35.57
Table 3: Performance comparison between VeOmni and TorchTitan with Qwen2-72B [72] on 128GPUs. Both configurations adopt mixed-precision training and full activation checkpointing, with a micro-batch size of 1 and a global batch size of 128.
Method Seqlen Memory (GB) Throughput MFU (%)
@@
FSDP+SP4 8k 58.62 54.08 565 499 28.04 24.74
FSDP+SP8 8k 58.35 51.73 330 297 16.11 14.75
FSDP+SP4 16k 62.15 64.35 624 544 35.34 30.44
FSDP+SP8 16k 59.98 58.02 532 394 29.82 22.03
FSDP+SP4 32k 68.35 79.36 592 497 41.05 34.16
FSDP+SP8 32k 63.49 70.39 533 398 36.72 27.31
FSDP+SP4 64k 79.10 OOM 468 - 43.98 -
FSDP+SP8 64k 70.09 79.04 441 335 41.63 31.56
FSDP+SP4 128k OOM OOM - - - -
FSDP+SP8 128k 76.67 OOM 303 - 43.68 -
Table 4: Performance of VeOmni with Qwen3-30B-A3B [58] on 128GPUs. We adopt mixed-precision training and full activation checkpointing, with a micro-batch size of 1 and a global batch size of 128. (TorchTitan does not support this setting.)
Method Seqlen Memory (GB) Throughput MFU (%)
FSDP+SP1+EP8 8k 45.58 2,776 7.45
FSDP+SP4+EP8 8k 22.56 1,216 3.02
FSDP+SP1+EP8 16k 69.99 2,853 10.74
FSDP+SP4+EP8 16k 34.32 1,897 6.55
FSDP+SP1+EP8 32k 77.25 2,417 13.40
FSDP+SP4+EP8 32k 63.08 2,064 11.35
FSDP+SP1+EP8 64k OOM - -
FSDP+SP4+EP8 64k 77.92 1,716 15.65
FSDP+SP1+EP8 128k OOM - -
FSDP+SP4+EP8 128k 78.01 1,075 17.92
These benchmark results demonstrate that VeOmni consistently achieves higher throughput and memory efficiency across a range of model sizes and sequence lengths. Notably, VeOmni supports long-sequence and MoE model training that exceed the memory or capability limits of TorchTitan.
Appendix B VeOmni’s API Design
B.1 Data Processing and Collating
Listing LABEL:lst:veomni-data-collator demonstrates the data collation pipeline in VeOmni, which provides a unified interface for processing and batching omni-modal inputs, including text, image, video, and audio. The OmniDataCollatorWithPacking class enables fine-grained control over feature packing and concatenation, making it straightforward to prepare batched training data for diverse modalities. This modular design simplifies the integration of new modalities and supports flexible token-level and feature-level alignment strategies. The resulting train_dataloader is constructed using build_dataloader(), which automatically aligns with the parallel training configuration.
Listing 1: Data Processing and Collating
@@ B.1 Data Processing and Collating
OmniDataCollatorWithPacking, build_dataloader
@@
B.2 Omni-Modal LLMs Initialization
Listing LABEL:lst:veomni-build-omni-model shows the usage of VeOmni’s build_omni_model API, which serves as the unified entry point for constructing omni-modal models. Given the configuration file and model component paths (including modality-specific encoders, decoders, and the foundation model), this function automatically assembles and initializes the model on a designated device. By abstracting away the underlying initialization logic, VeOmni ensures that users can easily instantiate complex omni-modal models in a modular and scalable fashion, while maintaining compatibility with distributed parallelization workflows.
Listing 2: Omni-Modal LLMs Initialization
@@ B.2 Omni-Modal LLMs Initialization
model = build_omni_model(
@@
B.3 Omni-Modal Model Computation
Listing LABEL:lst:veomni-train-omni-model-LABEL:lst:veomni-infer-omni-model illustrates the partial forward process implementation of the plug-and-play architecture in VeOmni, named OmniModel, which adopts an encoder–foundation–decoder architecture. The snippet demonstrates how the forward logic across training and inference modes integrates each module (i.e., encoder, foundation model, and decoder) in a seamless and modular pipeline. During inference, the control flow dynamically switches to modality-specific generation branches based on the generated text special tokens, enabling flexible omni-modal outputs. During inference, OmniModel supports parallelized generation and classifier-free guidance [68]. This architecture highlights the clear decoupling between encoder, foundation, and decoder modules, allowing each component to define and execute its own forward logic independently.
Listing 3: OmniModel Training Process
@@ B.3 Omni-Modal Model Computation
Listing 4: OmniModel Inference Process
class OmniModel:
@@
B.4 Parallel State Management
Listing LABEL:lst:veomni-parallel-state demonstrates the core interface of VeOmni’s parallel_state API, which provides a unified and declarative initialization of all n-D parallelism configurations. Unlike traditional methods [52] managing parallel states based on process groups, VeOmni leverages a DeviceMesh [36] abstraction to represent and organize various parallel topologies, including data parallelism (DP), tensor parallelism (TP), expert parallelism (EP), pipeline parallelism (PP), and sequence parallelism (SP). Once initialized via init_parallel_state, users can conveniently manage process groups, ranks, and communication topologies through the global parallel_state object. This design simplifies the orchestration of n-D parallelism strategies and decouples low-level distributed logic from model implementation.
Listing 5: Parallel State Management
from veomni.distributed.parallel_state import (
get_parallel_state, init_parallel_state